OpenStudy (anonymous):

Differentiate and simplify. y=(x+1)^4 x (x^2+1)^2?

OpenStudy (anonymous):

Would I use the product rule? and how would I write that out?

OpenStudy (anonymous):

\[9x^8+32x^7+56x^6+72x^5+70x^4+48x^3+24x^2+8x+1\]

OpenStudy (anonymous):

That's the answer? or is that the write out?

OpenStudy (anonymous):

answer according to me

OpenStudy (anonymous):

u meant to differentiate\[(x+1)^4*x*(x^2+1)^2\]

OpenStudy (anonymous):

if yes then this is the answer which i have given u

OpenStudy (anonymous):

Just (x+1)^4∗(x2+1)^2

OpenStudy (anonymous):

Would I just multiply them out 4 times? for the first parenthesis?

OpenStudy (anonymous):

then it is

OpenStudy (anonymous):

\[\huge 4(x+1)^3*(x^2+1)^2+4x(x+1)^4*(x^2+1)\]

OpenStudy (anonymous):

this hould be the answer!

OpenStudy (anonymous):

To be clear, differentiate means finding the derivative right? I'm loss on terms

OpenStudy (anonymous):

yes differentiation is finding the derivative

OpenStudy (anonymous):

apply product rule!!!

OpenStudy (anonymous):

How did you get 4x(x+1)^4?

OpenStudy (anonymous):

2.2x.(x+1)^4

OpenStudy (anonymous):

I don't get that =/

OpenStudy (anonymous):

Wow... Use chain rule and product rule. \[y=(x+1)^4(x^2+1)^2\] By product rule,\[\frac{d}{dx}[(x+1)^4(x^2+1)^2]=(x^2+1)^2\frac{d}{dx}(x+1)^4+(x+1)^4\frac{d}{dx}(x^2+1)^2\]By chain rule, i.e. \[\frac{dy(x)}{dx}=\frac{dy(u)}{du}\frac{du}{dx}\]We have\[\frac{d(x+1)^4}{x}=\frac{d(x+1)^4}{d(x+1)}\frac{d(x+1)}{dx}=4(x+1)^3\]and\[\frac{d(x^2+1)^2}{dx}=\frac{d(x^2+1)^2}{d(x^2+1)}\frac{d(x^2+1)}{dx}=2(x^2+1)(2x)=4x(x^2+1)\]Hence\[\frac{dy}{dx}=4(x^2+1)^2(x+1)^3+4x(x+1)^4(x^2+1)\]

OpenStudy (anonymous):

Woah. hold on o__O

OpenStudy (anonymous):

I don't see why the order has to be mixed up in the answer > <

OpenStudy (anonymous):

Just for better appearance and tidier.

OpenStudy (anonymous):

So could I write out 4(x+1)^3(x^2+1)?

OpenStudy (anonymous):

For factorization, you can do this.

OpenStudy (anonymous):

K